# Go Beyond Earth: Understanding Human Actions and Scenes in Microgravity Environments

### v1

@@ 5.2 Result Analysis for Fine-Grained Human Action Recognition
Table 1: Performance comparison of models fine-tuned on MicroG-4M, evaluated on the validation and test sets.
Model Validation Test Arch TC Backbone #Params (M) mAP (%) F1-score (%) Recall (%) AUROC (%) mAP (%) F1-score (%) Recall (%) AUROC (%) C2D [4] 8x8 R50 [77] $23.61$ $27.22$ $12.52$ $10.34$ $82.86$ $29.51$ $8.09$ $6.58$ $83.49$ C2D NLN [4] 8x8 R50 [77] $30.97$ $40.42$ $23.10$ $20.41$ $87.11$ $44.64$ $28.30$ $24.86$ $89.40$ I3D [62] 8x8 R50 [77] $27.33$ $40.93$ $19.78$ $16.93$ $86.44$ $46.41$ $26.37$ $22.25$ $88.79$ I3D NLN [62] 8x8 R50 [77] $34.68$ $41.42$ $24.11$ $23.00$ $86.37$ $47.12$ $28.07$ $24.65$ $88.52$ Slow [4] 8x8 R50 [77] $31.74$ $40.32$ $21.83$ $19.08$ $84.55$ $45.19$ $26.13$ $22.77$ $88.49$ Slow [4] 4x16 R50 [77] $31.74$ $42.97$ $22.73$ $19.71$ $85.46$ $46.37$ $28.72$ $25.38$ $88.30$ SlowFast [4] 8x8 R50 [77] $33.76$ $38.76$ $20.29$ $17.66$ $85.91$ $43.02$ $22.63$ $18.98$ $88.51$ SlowFast [4] 4x16 R50 [77] $33.76$ $37.10$ $17.74$ $14.90$ $84.94$ $42.10$ $23.69$ $20.18$ $87.54$ MViTv1 [61] 16x4 B-CONV $36.34$ $17.79$ $7.89$ $6.86$ $72.40$ $12.86$ $5.54$ $4.66$ $74.63$ MViTv2 [16] 16x4 S $34.27$ $17.57$ $8.31$ $6.92$ $72.67$ $15.14$ $8.16$ $7.17$ $78.61$ X3D [3] 13x6 S $2.02$ $17.59$ $6.63$ $5.63$ $78.27$ $14.07$ $5.77$ $4.52$ $78.23$ X3D [3] 16x5 L $4.37$ $23.56$ $8.82$ $7.38$ $80.56$ $18.70$ $9.15$ $7.47$ $78.27$
@@
Note: All models has been pretrained on Kinetics400 [11] dataset and continually trained on MicroG-4M. TC denotes the temporal configuration (frame length × sampling rate). #Params indicates the number of parameters (in millions, M).
@@
Table 2: Zero-shot performance on MicroG-4M test set for models pretrained on Kinetics and fine-tuned on AVA [12].
@@
Model Test Result Arch TC Backbone Pretrain Fine-tune mAP (%) F1-score (%) Recall (%) AUROC (%) Slow [4] 8x8 R50 [77] Kinetics 400 [11] AVA v2.2 [12] $16.24$ $2.67$ $1.99$ $73.83$ SlowFast [4] 32x2 R101 [77] Kinetics 600 [78] AVA v2.2 [12] $23.81$ $6.32$ $6.62$ $77.83$
@@
Note: All metrics are macro-averaged over action classes. mAP is measured at IoU $=0.5$. F1 and AUROC are computed per class and then averaged. TC denotes the temporal configuration (frame length × sampling rate).
@@
Figure 3: Qualitative results for fine-grained human action recognition in microgravity, where GT denotes ground truth, MicroG-4M indicates predictions from Slow [4] finetuned on MicroG-4M, and AVA denotes predictions from the same model finetuned on AVA [12]. The MicroG-4M model provides more accurate predictions than its Earth-trained counterpart.
We evaluate several representative models on MicroG-4M, all pretrained on Kinetics400 [11] and fine-tuned on our dataset. As shown in Table 1, most models plateau around $47\%$ test mAP, highlighting a substantial domain gap between Earth-based datasets and microgravity. The ranking of models deviates from trends on Kinetics400 [11] or AVA [12], e.g., convolutional models with non-local enhancements [4, 62] achieve the highest mAP and AUROC, while Slow (4×16) [4] yields the best F1-score, outperforming newer transformer-based designs [3, 16]. This suggests local spatial modeling and structured receptive fields remain advantageous in microgravity, where motion lacks gravitational consistency. Moreover, models with longer temporal windows perform better, reinforcing the importance of domain-adapted temporal reasoning for capturing slow or non-linear microgravity dynamics.
Comparing to AVA-finetuned models (Table 2), we find a substantial drop in performance when transferring directly from AVA [12] to MicroG-4M, even when pretraining and backbone are kept constant. For example, SlowFast trained on AVA [12] achieves only $23.81\%$ in mAP on MicroG-4M, significantly lower than its performance when fine-tuned on our dataset. This reinforces that Earth-trained models carry strong but fragile assumptions about vertical orientation, motion causality, and object affordances, which no longer hold in microgravity.
Together, these findings validate MicroG-4M’s role as a diagnostic benchmark that breaks gravity-dependent priors and challenges models to reason over irregular motion, spatial ambiguity, and contact-free interactions. Our dataset thus provides a rigorous testbed for evaluating and advancing space-adapted video understanding models, especially in the context of astronaut assistance and autonomous system development.
Qualitative Analysis. Figure 3 presents qualitative comparisons between models trained on AVA [12] and those fine-tuned on MicroG-4M, across 5 representative video clips captured inside and outside spacecraft cabins. The MicroG-4M–trained model demonstrates high alignment with ground truth labels for core actions such as “Stand”, “Walk”, and “Talk-to”, while the AVA [12]-based counterpart consistently misinterprets floating or inverted postures as “Bend/Bow” or “Sit”, revealing its reliance on Earth-centric gravitational priors.
@@
A representative example in the fifth column shows the AVA [12]-finetuned model misclassifying a floating astronaut as “Sit” while the MicroG-4M–trained model correctly predicts “Stand” reflecting a common correction of gravity-induced biases. Models trained on MicroG-4M also demonstrate improved robustness in distinguishing passive object drift from intentional manipulation, though semantic ambiguity remains, e.g., predicting “Carry/Hold” when a tool drifts near the astronaut’s hand without actual interaction. These results indicate that MicroG-4M mitigates terrestrial biases, enhances sensitivity to body-object dynamics, and better captures domain-specific actions, supporting future work on temporal coherence and intentionality modeling.
@@ 5.3 Evaluation of Video Captioning Models
Open Sourced
Video-ChatGPT [79] 3 0.06 0.12 10.10 4.33 39.61 85.40
mPLUG-Owl3 [64] 3 0.16 0.40 11.87 5.88 47.45 85.91
LLaVA-NeXT [80] 8 0.30 1.88 16.32 14.45 54.16 84.98
Video-LLaVA [66] 8 0.03 0.07 9.29 4.12 42.97 84.89
Qwen2.5-VL [81] 9 0.03 1.34 13.75 15.67 56.46 84.01
InternVideo [82] 90 0.77 2.60 16.57 15.18 55.28 85.41
Closed Sourced
GPT-4o [19] 6 1.74 2.65 16.46 11.27 62.18 86.75
Gemini 1.5 Pro [68] 16 3.52 3.28 17.34 15.19 63.38 86.25
The results in Table 3 reveal how video-language models perform on the MicroG-4M benchmark, highlighting key challenges introduced by microgravity-specific content. Lexical metrics such as CIDEr and BLEU-4 show low overall values, especially among open-source models, suggesting a significant distributional shift between MicroG-4M and typical pretraining data. The dataset’s domain-specific vocabulary, visually compositional scenes, and semantically dense annotations likely reduce surface-level overlap, which these metrics are sensitive to. In contrast, semantic similarity metrics such as S-BERT and BERTScore remain relatively higher and more consistent, indicating that several models capture the underlying intent even without lexical alignment. This underscores the semantic richness of MicroG-4M, where alternative phrasings and scientific terminology often convey similar meanings. Performance differences further reveal that input frame density and pretraining modality play key roles. InternVideo [17], which processes $90$ frames sampled within a $3$s window, consistently outperforms other open-source models. This suggests that dense sampling, coupled with video-specific pretraining, enhances the model’s ability to capture subtle spatial patterns and object-scene relationships—features that are particularly important in microgravity scenarios, where visual cues are often atypical or physically ambiguous. Closed-source models, i.e., GPT-4o [19] and Gemini 1.5 Pro [68] achieve better scores on both lexical and semantic metrics, likely due to broader data exposure, larger capacity, or more advanced cross-modal fusion strategies. However, their relatively small performance gains further validate the challenge posed by MicroG-4M. In general, these findings position MicroG-4M as a demanding benchmark for evaluating multimodal models under domain change, highlighting the need for robust spatial reasoning, domain adaptation, and semantically aware generation strategies in unconventional environments.
@@ 5.4 Evaluation of Visual Question Answering Models
Table 4: Performance comparison of Visual Question Answering (VQA) models on the MicroG-4M benchmark. “#f” denotes the number of input frames used during model inference.
@@
Open Sourced
LLaVA-NeXT [80] 8 24.00 22.14 15.56 12.40 38.08 87.15
Video-LLaVA [66] 8 25.70 28.47 15.90 10.71 35.39 87.13
Qwen2.5-VL [81] 9 2.99 0.65 8.35 8.47 40.65 84.80
Closed Sourced
Gemini 1.5 Pro [68] 16 8.78 1.33 13.03 12.54 43.15 86.41
GPT-4o [19] 6 33.98 3.76 18.11 15.89 44.56 87.81
The results in Table 4 demonstrate that MicroG-4M presents distinct challenges for visual question answering. Notably, there is a significant divergence between lexical and semantic evaluation metrics, particularly among open-source models. For example, Qwen2.5-VL [67] yields a BLEU-4 of only 0.65 and a CIDEr score of 2.99, yet achieves the highest S-VQA score in its category (40.65). This contrast suggests that MicroG-4M questions often admit multiple semantically valid answers that differ lexically, such as paraphrased actions, scientific terms, or object references adapted to microgravity settings. This characteristic does not reflect inconsistency in evaluation, but rather underscores the conceptual and linguistic diversity embedded in the dataset. In addition, the moderate absolute scores of even the top-performing closed-source models, such as GPT-4o [19] (CIDEr 33.98, S-VQA 44.56), reveal the difficulty of reasoning over visual content in microgravity. Unlike conventional VQA datasets, MicroG-4M includes visually ambiguous cues, e.g., floating objects, unusual body orientations, and tool manipulations under microgravity that challenge models trained primarily on terrestrial data. This suggests that current pretraining corpora lack sufficient coverage of such scenarios, and that purely scaling model capacity is insufficient for reliable generalization.
Interestingly, increasing the number of input frames within the fixed $3$s window does not consistently yield better performance. For example, Gemini 1.5 Pro [68] processes 16 frames but performs worse than GPT-4o [19], which uses only 6 frames. This indicates that dense frame sampling alone is insufficient. Instead, performance depends more critically on the model’s ability to extract semantically salient cues, e.g., astronaut posture, object manipulation, and spatial configurations, from visually subtle or low-motion segments. In microgravity environments, where conventional motion dynamics and object affordances are altered, effective spatial reasoning and cross-modal alignment appear to be more decisive than temporal redundancy. In summary, MicroG-4M reveals key limitations of current VQA systems in addressing domain-specific challenges, particularly those involving spatial complexity, ambiguous motion, and semantically flexible queries inherent to microgravity. Its comprehensive and specialized design establishes it as a valuable testbed for probing the robustness, adaptability, and generalization capabilities of multimodal models well beyond the scope of conventional Earth-based benchmarks.
@@ 6 Conclusion
In this work, we present MicroG-4M, the first large-scale dataset specifically curated for human action recognition and vision-language understanding in microgravity environments. The dataset features $4,759$ annotated video clips with over $390,000$ bounding boxes and $13,000$+ action labels across $50$ unique action classes. It also includes human-written captions and over $7,000$ VQA pairs, enabling rich semantic understanding and reasoning. To facilitate fair evaluation, we introduce MicroG-Bench, a benchmark that systematically assesses state-of-the-art models across fine-grained action recognition, video captioning, and question answering tasks. Our results show that existing deep learning models degrade significantly in space-like settings, emphasizing the need for dedicated benchmarks and domain-specific adaptation. MicroG-4M aims to advance robust, generalizable AI for astronaut support and autonomous space operations.
Limitations: The MicroG-4M dataset is currently limited to RGB input and includes both real and simulated footage, introducing potential modality and domain biases. Additionally, while the annotated subset is suitable for benchmarking, the scale of the dataset can be further enlarged to enable more space-related exploration for the human perception models. More illustrations can be found in the supplementary materials.
ACKNOWLEDGMENT

### v3

@@ 5.2 Result Analysis for Fine-Grained Human Action Recognition
Table 1: Performance of models fine-tuned on MicroG-4M, evaluated on the validation and test sets.
[…]
@@
Note: All models have been pretrained on Kinetics400 Kay et al. (2017) dataset and continually trained on MicroG-4M. TC denotes the temporal configuration (frame length × sampling rate). #Params indicates the number of parameters (in millions, M).
@@
Table 2: Zero-shot performance on MicroG-4M test set for models pretrained on Kinetics and fine-tuned on AVA Gu et al. (2018).
@@
Model Test Result Arch TC Backbone Pretrain Fine-tune mAP (%) F1-score (%) Recall (%) AUROC (%) Slow Feichtenhofer et al. (2019) 8x8 R50 He et al. (2016) Kinetics 400 Kay et al. (2017) AVA v2.2 Gu et al. (2018) $16.24$ $2.67$ $1.99$ $73.83$ SlowFast Feichtenhofer et al. (2019) 32x2 R101 He et al. (2016) Kinetics 600 Carreira et al. (2018) AVA v2.2 Gu et al. (2018) $23.81$ $6.32$ $6.62$ $77.83$
@@
Note: All metrics are macro-averaged over action classes. mAP is measured at IoU =0.5=0.5. F1 and AUROC are computed per class and then averaged. TC denotes the temporal configuration (frame length × sampling rate).
Table 3: Cross-domain transfer under matched AVA fine-tuning (zero-shot evaluation).
Model TC Backbone Test Set mAP (%) AUROC (%)
SlowFast Feichtenhofer et al. (2019) $32{\times}2$ R101 He et al. (2016) JHMDB Jhuang et al. (2013) 47.50 83.98
SlowFast Feichtenhofer et al. (2019) $32{\times}2$ R101 He et al. (2016) MicroG-4M 23.81 77.83
Slow Feichtenhofer et al. (2019) $8{\times}8$ R50 He et al. (2016) JHMDB Jhuang et al. (2013) 34.24 76.96
Slow Feichtenhofer et al. (2019) $8{\times}8$ R50 He et al. (2016) MicroG-4M 16.24 73.83
@@
Figure 3: Qualitative results for fine-grained human action recognition in microgravity, where GT denotes ground truth, MicroG-4M indicates predictions from Slow fine-tuned on MicroG-4M, and AVA denotes predictions from the same model fine-tuned on AVA. The MicroG-4M model provides more accurate predictions than its Earth-trained counterpart.
[…]
@@
Qualitative Analysis. Figure 3 presents qualitative comparisons between models trained on AVA Gu et al. (2018) and those fine-tuned on MicroG-4M, across 5 representative video clips captured inside and outside spacecraft cabins. The MicroG-4M–trained model demonstrates high alignment with ground truth labels for core actions such as “Stand”, “Walk”, and “Talk-to”, while the AVA Gu et al. (2018)-based counterpart consistently misinterprets floating or inverted postures as “Bend/Bow” or “Sit”, revealing its reliance on Earth-centric gravitational priors. A representative example in the fifth column shows the AVA Gu et al. (2018)-finetuned model misclassifying a floating astronaut as “Sit” while the MicroG-4M–trained model correctly predicts “Stand” reflecting a common correction of gravity-induced biases. Models trained on MicroG-4M also demonstrate improved robustness in distinguishing passive object drift from intentional manipulation, though semantic ambiguity remains, e.g., predicting “Carry/Hold” when a tool drifts near the astronaut’s hand without actual interaction. These results indicate that MicroG-4M mitigates terrestrial biases, enhances sensitivity to body-object dynamics, and better captures domain-specific actions, supporting future work on temporal coherence and intentionality modeling.
@@ 5.3 Evaluation of Video Captioning Models
[…]
Video-ChatGPT Maaz et al. (2024) 3 0.06 0.12 10.10 4.33 39.61 85.40
mPLUG-Owl3 Ye et al. (2025) 3 0.16 0.40 11.87 5.88 47.45 85.91
LLaVA-NeXT Zhang et al. (2024) 8 0.30 1.88 16.32 14.45 54.16 84.98
Video-LLaVA Lin et al. (2024) 8 0.03 0.07 9.29 4.12 42.97 84.89
Qwen2.5-VL Qwen et al. (2025) 9 0.03 1.34 13.75 15.67 56.46 84.01
Tarsier2-Recap-7B Yuan et al. (2025) 16 0.04 0.03 0.17 0.12 51.35 84.53
InternVideo Xing et al. (2024) 90 0.77 2.60 16.57 15.18 55.28 85.41
[…]
GPT-4o Hurst et al. (2024) 6 1.74 2.65 16.46 11.27 62.18 86.75
Gemini 1.5 Pro Reid et al. (2024) 16 3.52 3.28 17.34 15.19 63.38 86.25
The results in Table 5.3 reveal how video-language models perform on the MicroG-4M benchmark, highlighting key challenges introduced by microgravity-specific content. Lexical metrics such as CIDEr and BLEU-4 show low overall values, especially among open-source models, suggesting a significant distributional shift between MicroG-4M and typical pretraining data. The dataset’s domain-specific vocabulary, visually compositional scenes, and semantically dense annotations likely reduce surface-level overlap, which these metrics are sensitive to. In contrast, semantic similarity metrics such as S-BERT and BERTScore remain relatively higher and more consistent, indicating that several models capture the underlying intent even without lexical alignment. This underscores the semantic richness of MicroG-4M, where alternative phrasings and scientific terminology often convey similar meanings. Performance differences further reveal that input frame density and pretraining modality play key roles. InternVideo Wang et al. (2022), which processes $90$ frames sampled within a $3$s window, consistently outperforms other open-source models. This suggests that dense sampling, coupled with video-specific pretraining, enhances the model’s ability to capture subtle spatial patterns and object-scene relationships—features that are particularly important in microgravity scenarios, where visual cues are often atypical or physically ambiguous. Closed-source models, i.e., GPT-4o Hurst et al. (2024) and Gemini 1.5 Pro Reid et al. (2024) achieve better scores on both lexical and semantic metrics, likely due to broader data exposure, larger capacity, or more advanced cross-modal fusion strategies. However, their relatively small performance gains further validate the challenge posed by MicroG-4M. In general, these findings position MicroG-4M as a demanding benchmark for evaluating multimodal models under domain change, highlighting the need for robust spatial reasoning, domain adaptation, and semantically aware generation strategies in unconventional environments.
@@ 5.4 Evaluation of Visual Question Answering Models
Table 5: Experiments of Visual Question Answering (VQA) models on the MicroG-4M benchmark. “#f” denotes the number of input frames used during model inference.
@@
[…]
LLaVA-NeXT Zhang et al. (2024) 8 24.00 22.14 15.56 12.40 38.08 87.15
Video-LLaVA Lin et al. (2024) 8 25.70 28.47 15.90 10.71 35.39 87.13
Qwen2.5-VL Qwen et al. (2025) 9 2.99 0.65 8.35 8.47 40.65 84.80
Tarsier2-Recap-7B Yuan et al. (2025) 16 5.08 0.01 0.08 0.09 29.60 85.30
[…]
Gemini 1.5 Pro Reid et al. (2024) 16 8.78 1.33 13.03 12.54 43.15 86.41
GPT-4o Hurst et al. (2024) 6 33.98 3.76 18.11 15.89 44.56 87.81
The results in Table 5.4 demonstrate that MicroG-4M presents distinct challenges for visual question answering. Notably, there is a significant divergence between lexical and semantic evaluation metrics, particularly among open-source models. For example, Qwen2.5-VL Bai et al. (2025) yields a BLEU-4 of only 0.65 and a CIDEr score of 2.99, yet achieves the highest S-VQA score in its category (40.65). This contrast suggests that MicroG-4M questions often admit multiple semantically valid answers that differ lexically, such as paraphrased actions, scientific terms, or object references adapted to microgravity settings. This characteristic does not reflect inconsistency in evaluation, but rather underscores the conceptual and linguistic diversity embedded in the dataset. In addition, the moderate absolute scores of even the top-performing closed-source models, such as GPT-4o Hurst et al. (2024) (CIDEr 33.98, S-VQA 44.56), reveal the difficulty of reasoning over visual content in microgravity. Unlike conventional VQA datasets, MicroG-4M includes visually ambiguous cues, e.g., floating objects, unusual body orientations, and tool manipulations under microgravity that challenge models trained primarily on terrestrial data. This suggests that current pretraining corpora lack sufficient coverage of such scenarios, and that purely scaling model capacity is insufficient for reliable generalization. Interestingly, increasing the number of input frames within the fixed $3$s window does not consistently yield better performance. For example, Gemini 1.5 Pro Reid et al. (2024) processes 16 frames but performs worse than GPT-4o Hurst et al. (2024), which uses only 6 frames. This indicates that dense frame sampling alone is insufficient. Instead, performance depends more critically on the model’s ability to extract semantically salient cues, e.g., astronaut posture, object manipulation, and spatial configurations, from visually subtle or low-motion segments. In microgravity environments, where conventional motion dynamics and object affordances are altered, effective spatial reasoning and cross-modal alignment appear to be more decisive than temporal redundancy. In summary, MicroG-4M reveals key limitations of current VQA systems in addressing domain-specific challenges, particularly those involving spatial complexity, ambiguous motion, and semantically flexible queries inherent to microgravity. Its comprehensive and specialized design establishes it as a valuable testbed for probing the robustness, adaptability, and generalization capabilities of multimodal models well beyond the scope of conventional Earth-based benchmarks.
@@ 6 Conclusion
In this work, we present MicroG-4M, the first large-scale dataset specifically curated for human action recognition and vision-language understanding in microgravity environments. The dataset features $4,759$ annotated video clips with over $390,000$ bounding boxes and 13,00013,000+ action labels across $50$ unique action classes. It also includes human-written captions and over $7,000$ VQA pairs, enabling rich semantic understanding and reasoning. We introduce MicroG-Bench, a benchmark for evaluating state-of-the-art models in fine-grained action recognition, video captioning, and question answering. Results show significant performance degradation in space-like settings, highlighting the need for domain-specific benchmarks and adaptation. MicroG-4M advances robust, generalizable AI for astronaut support and autonomous space operations.
Acknowledgement